A 0.1 kg steel ball falls from a height of 10m on the ground and rebounds to a height of 7m. If the dissipated energy remains in the ball in the form of heat then by how much will the temperature of the ball be raised

Grade:11

1 Answers

Akash
5 years ago
Mgh=ms[T2-T1]
Gh=s[T2-T1]
Gh=9.8*3=29.4J
S=110 cal
1cal=4.2J
110cal=462J
[T2-T1]=29.4/462
[T2-T1]=0.064C
 
110cal=110*4.2J
110cal=